OpenStudy (anonymous):

Which of the following could be the equation of the graph shown below? Check all that apply.

OpenStudy (♪chibiterasu):

We need to see the graph and the options :/

OpenStudy (anonymous):

A. x - y = 9 B. y = -2x + 5 C. y = 5 D. 2x + y = 4